# A Bayesian latent Gaussian conditional autoregressive copula model for analyzing spatially-varying trends in rainfall

### v1

@@ Abstract
Assessing the availability of rainfall water plays a crucial role in rainfed agriculture. Given the substantial proportion of agricultural practices in India being rainfed and considering the potential trends in rainfall amounts across years due to climate change, we build a statistical model for analyzing annual total rainfall data for 34 meteorological subdivisions of mainland India available for 1951–2014. Here, we model the margins using a gamma regression model and the dependence through a Gaussian conditional autoregressive (CAR) copula model. Due to the natural variation of the average annual rainfall received across various dry through wet regions of the country, we allow areally-varying gamma regression coefficients under a latent Gaussian model framework. The neighborhood structure of the regions determines the dependence structure of both the likelihood and the prior layers, where we explore both CAR and intrinsic CAR structures for the priors. The proposed methodology also imputes the missing data effectively. We use the Markov chain Monte Carlo algorithms to draw Bayesian inferences. In simulation studies, the proposed model outperforms some competitors that do not allow a dependence structure at the data or prior layers. Implementing the proposed method for the Indian areal rainfall dataset, we draw inferences about the model parameters and discuss the potential effect of climate change on rainfall across India.
@@ 1 Introduction
Rainfall is a critical climatic factor influencing various sectors, including agriculture, hydrology, and disaster management. It is essential to analyze and model the spatial and temporal variations in rainfall, especially in monsoon-driven regions like India; with more than 60% of its agricultural land reliant on rainfed farming, India has the most considerable such extent globally [1, 2]. Climate change has intensified these uncertainties in recent decades, highlighting the need for statistical models that effectively capture rainfall patterns and dependencies [3]. Climate change alters weather patterns, and these shifts influence the distribution of temperature, wind, and rainfall, leading to variations in spatial rainfall patterns. While some regions may experience increased and more frequent rainfall, others may face droughts and declining precipitation. With varying marginal behavior across different global regions, rainfall exhibits inherent spatial dependence, shaped by large-scale atmospheric circulation, geographical features, and local climatic conditions [4, 5]. Hence, these phenomena motivate us to study varying trends in rainfall across the meteorological subdivisions of India while allowing a realistic spatial dependence structure exhibited by the data.
@@
Historically, the statistical modeling of monthly, seasonal, or annual total rainfall has remained an important research area in meteorology. Annual total rainfall data are usually nonzero, and the histograms appear positively skewed; for all meteorological subdivisions of India, the recorded annual total rainfall data do not include any year with entirely nil rainfall. Hence, the justified probability distributions here are right-skewed and supported over the positive real line; some examples are exponential [6, 7], gamma [8, 9], log-normal [10, 11], Weibull [12, 13], and generalized exponential [14, 15] distributions. However, meteorologists have historically favored the gamma distribution for modeling rainfall data [16]. While the data across years are often assumed to be distributed identically in the meteorology literature, it is more appropriate to consider a nonstationary behavior across years in the marginal distribution due to the potential trends in meteorological variables driven by climate change. For this purpose, a gamma regression model is a possible approach that various researchers have explored for different scientific disciplines [17, 18].
@@
While the gamma regression can effectively model the nonstationary marginal behavior in rainfall, a fundamental limitation of many traditional univariate models is their failure to incorporate spatial dependency, which is critical for accurate estimation and prediction [19]. Copulas [20, 21, 22] are powerful statistical tools that enable the modeling of complex dependencies between multiple random variables, even extending beyond traditional linear correlation measures [23, 24]. In spatial statistics, copulas facilitate the construction of joint distributions by separately modeling marginal distributions and their spatial dependence structure [25, 26]; this separation allows for greater flexibility in capturing nonlinear and asymmetric dependencies that are often present in spatial data, leading to improved inference and prediction in various applications [27, 28, 29, 30]. [31] attempt to combine the copula theory with the entropy theory for bivariate rainfall and runoff analysis. Some other applications of copulas, e.g., Archimedean copulas for meteorological data, are covered in [32, 33]. [34] predict temporal trends of precipitation and temperature using copula under a climate change scenario. In our context of areal data, [35] propose a copula-based hierarchical model with covariance selection for unbiased estimation of marginal parameters, providing a dependence structure with intuitive conditional and marginal interpretations. They develop a computational framework that permits efficient frequentist inference for their model, even for large datasets.
@@
Several studies have demonstrated the efficacy of using gamma marginal distributions combined with copula models to analyze rainfall patterns. For instance, [36] employ gamma distributions as marginals within different copula models to capture spatial dependencies in precipitation data. On the other hand, [37] introduces a simulation-based method to enhance the preservation of cross-correlations in multisite precipitation simulations, addressing limitations observed in direct and indirect estimation methods. Here, the author explores a Gaussian copula with identical gamma marginal distributions for data collected at twelve nearby spatial locations; however, the model assumptions are unsuitable for a large geographical domain. Similarly, [38] utilize copulas with gamma marginal distributions to model the relationship between coarse- and fine-scale rainfall depths. The study in [39] describes precipitation intensity using space and time-dependent gamma distributions. While most of these researches do not assume a gamma regression framework for the margins, the rest do not focus on annual trends. Besides, the copulas used in these papers are for continuous spatial processes; hence, they are not directly applicable to areal data modeling.
Outside the avenue of copula-based modeling, the two most popular frameworks for areal data modeling are conditional autoregressive (CAR) models [40] and simultaneous autoregressive (SAR) models [41]. While SAR models are used widely in econometrics, regional science, and social sciences, CAR models are popular in Bayesian spatial statistics, epidemiology, and environmental studies. Among numerous uses of the CAR model in the environmental statistics literature, [42] use hierarchical CAR models for temperature variability analysis, [43] analyze flood risks and rainfall patterns using CAR models, and [44] use integrated CAR models with spatial dynamic processes for environmental prediction. In most of these models, the data are assumed to be Gaussian, conditioning on the parameters and the latent variables. However, using the CAR model as a copula in the context of environmental statistics or other scientific disciplines is not known.
@@
Given the large spatial extent of the meteorological subdivisions of India, with some subdivisions in the western parts of the country receiving very low rainfall and some subdivisions in the east covering the wettest locations of the globe, the requirement of areally-varying gamma regression coefficients is obvious [45]. Besides, the regression coefficients are likely to express local homogeneity, and hence, we can model the priors for the areally-varying coefficients under the framework of Gaussian graphical models [42]. The CAR and intrinsic CAR [ICAR, 46] priors are popular choices for areally-varying coefficients. While [44] discuss ICAR priors in spatio-temporal ecological models, [42] use ICAR priors for modeling spatial rainfall variability. [47], [48], and many others use the CAR priors, while the R package CARBayes [49] implements some Bayesian areal data models with CAR priors.
@@
In this paper, we develop a statistical model to analyze annual total rainfall data for 34 meteorological subdivisions of mainland India from 1951 to 2014. Given the suitability of the gamma distribution for modeling rainfall data and the possible trend in the marginal distribution due to climate change, we assume a gamma regression model for the marginal distributions, where we consider a logarithmic link for the mean structure, written as a linear function of the years (suitably centered and scaled). We consider a Gaussian conditional autoregressive (CAR) copula for modeling the dependence structure where the adjacency is determined by whether or not two meteorological subdivisions share a boundary. While the usual CAR models show nonstationary marginal distributions due to the boundary effects, we appropriately scale the CAR covariance matrix to ensure the copula conditions are satisfied. We allow varying gamma regression coefficients within a latent Gaussian model framework to account for the natural variation in average annual rainfall across different climatic regions. Similar to the likelihood layer, the neighborhood structure of the regions determines the dependence structure in prior layers, where we explore CAR and ICAR priors. Our methodology also effectively imputes missing data. We draw Bayesian inference using Markov chain Monte Carlo algorithms, specifically, using a Metropolis-within-Gibbs sampler. To explore the benefit of considering the dependence structures at the likelihood and prior layers, we conduct an extensive simulation study by simulating many datasets of the same dimension as the Indian subdivision-wise rainfall dataset after assuming moderate through high spatial correlation. Using various criteria for model comparison, we then compare the performances of the proposed and competing models in terms of model fitting and estimation uncertainties. Applying this method to the Indian areal rainfall dataset, we draw inferences about the model parameters and examine the potential impacts of climate change on rainfall patterns across India.
@@
We structure the remainder of this article as follows. Section 2 provides a brief background on gamma regression, conditional autoregressive models, and copulas. In Section 3, we describe the details of the Indian rainfall dataset we analyze and provide an exploratory analysis motivating us to choose specific components in the proposed statistical methodology. Section 4 details the construction of the conditional autoregressive copula for modeling spatial dependence, a gamma regression with areally-varying coefficients for margins, and the CAR and ICAR prior structures for the coefficients. In Section 5, we discuss an MCMC sampling strategy employed for parameter estimation. Section 6 discusses a simulation study to validate the correctness of the proposed parameter estimation procedure and to compare the proposed model with some simpler alternatives that ignore spatial dependence structures. In Section 7, we apply the proposed and competing methods to the Indian rainfall dataset and discuss the results. Finally, Section 8 concludes with a discussion on key findings, implications for climate studies, and future research directions.
@@ 2 Background
This section briefly overviews some key statistical tools required for our methodology and data analysis, including gamma regression for handling nonstationary skewed positive-valued data, conditional autoregressive (CAR) models for areal datasets, and copula. We also discuss two exploratory tools, including the Quantile-Quantile (QQ) plot of uniform-transformed data for marginal model validation and Moran’s I for spatial autocorrelation assessment. Finally, we mention the Deviance information criterion (DIC) and the Watanabe-Akaike information criterion (WAIC) for Bayesian model comparison.
@@ 2.1 Gamma regression
In a gamma regression model for $n$ independent but non-identically distributed positive-valued observations, we assume that the response variable $Y_{i}$ follows
[…]
@@
$\lambda_{i}=a/\mu_{i}=a~{}\exp[-\bm{x}_{i}^{\top}\bm{\beta}]=\exp[\log(a)-\bm{x}_{i}^{\top}\bm{\beta}].$
Here, the coefficient of variation $\sqrt{\text{Var}(Y_{i})}/\text{E}(Y_{i})=a^{-1/2}$ does not depend on $\mu_{i}$.
2.2 Conditional Autoregressive (CAR) model
The CAR model assumes that the conditional distribution of the response variable at each region, given all other areas, depends only on the responses of the neighboring areas. Given a set of regionally indexed observations, the CAR model specifies a (zero-mean) Gaussian Markov Random Field (GMRF) structure
$Y_{i}|Y_{j},j\neq i\sim\textrm{Normal}\left(\rho\sum_{j\sim i}w_{ij}Y_{j},\frac{\sigma^{2}}{m_{i}}\right),$ (3)
where $w_{ij}$ are spatial weights defining the neighborhood structure, $\rho$ is the spatial autocorrelation parameter, and $m_{i}=\sum_{j\sim i}w_{ij}$. Using Brook’s lemma [50], the covariance matrix for all regions is given by $\sigma^{2}(\bm{M}-\rho\bm{W})^{-1},$ where $\bm{W}$ is the spatial weight matrix. The most common choice of $w_{ij}$s, the $(i,j)$th element of $\bm{W}$, is
$w_{ij}=\begin{cases}1~{}~{}\text{ if }~{}~{}i\sim j\\ 0~{}~{}\text{ if }~{}~{}i\nsim j\end{cases},$ (4)
where $i\sim j$ indicates regions $i$ and $j$ are neighbors. Here, $i\nsim i$, that is $w_{ii}=0$ for all $i$. The correlation parameter $\rho$ controls the strength of dependence among the regions. Here, $\bm{M}$ is a diagonal matrix with its $(i,i)$th element $m_{ii}=\sum_{j}w_{ij},$ and $\sigma^{2}$ is a marginal variance parameter that controls the overall spatial covariance. The CAR models enable local spatial smoothing, ensuring that geographically close observations share information, thereby improving predictive accuracy and inference. Multiple types of CAR models are considered based on the situation. A proper CAR model is given in (3). When $\rho=1$, the corresponding CAR model is called an intrinsic CAR (ICAR) model. The covariance matrix is singular for an ICAR model and hence, it can only be used as improper priors for areally-varying coefficients but not for modeling the likelihood.

### v2

@@ Abstract
Assessing the availability of rainfall water plays a crucial role in rainfed agriculture. Given the substantial proportion of agricultural practices in India being rainfed and considering the potential trends in rainfall amounts across years due to climate change, we build a statistical model for analyzing monsoon total rainfall data for 34 meteorological subdivisions of mainland India available for 1951–2014. Here, we model the marginal distributions using a gamma regression model and the dependence through a Gaussian conditional autoregressive (CAR) copula model. Due to the natural variation in the monsoon total rainfall received across various dry through wet regions of the country, we allow the parameters of the marginal distributions to be spatially varying, under a latent Gaussian model framework. The neighborhood structure of the regions determines the dependence structure of both the likelihood and the prior layers, where we explore both CAR and intrinsic CAR structures for the priors. The proposed methodology also effectively imputes the missing data. We use the Markov chain Monte Carlo algorithms to draw Bayesian inferences. In simulation studies, the proposed model outperforms several competitors that do not allow a dependence structure at the data or prior layers. Implementing the proposed method for the Indian areal rainfall dataset, we draw inferences about the model parameters and discuss the potential effect of climate change on rainfall across India. While the assessment of the impact of climate change on rainfall motivates our study, the proposed methodology can be easily adapted to other contexts dealing with non-Gaussian non-stationary areal datasets where data from single or multiple temporal covariates are also available, and it is appropriate to assume their coefficients to be spatially varying.
@@ 1 Introduction
Rainfall is a crucial climatic factor that influences various sectors, including agriculture, hydrology, and disaster management. It is essential to analyze and model the spatial and temporal variations in rainfall, especially in monsoon-driven regions like India, where more than 60% of its agricultural land relies on rainfed farming, making India the most considerable such region globally [venkateswarlu2011rainfed, hazra2024robust]. Climate change has intensified these uncertainties in recent decades, highlighting the need for statistical models that effectively capture rainfall patterns and dependencies [allan2023intergovernmental]. Climate change alters weather patterns, and these shifts influence the distribution of temperature, wind, and rainfall, leading to variations in spatial rainfall patterns. While certain regions may experience increased and more frequent rainfall, others may face droughts and declining precipitation. With varying marginal behavior across different global regions, rainfall exhibits inherent spatial dependence, shaped by large-scale atmospheric circulation, geographical features, and local climatic conditions [guhathakurta2008trends, hazra2024efficient]. Hence, these phenomena motivate us to study the varying trends in rainfall across the meteorological subdivisions of India, while allowing for a realistic spatial dependence structure exhibited by the data.
@@
Historically, the statistical modeling of monthly, seasonal, or annual total rainfall has remained an important research area in meteorology. Monsoon total rainfall data are usually nonzero, and the histograms appear positively skewed; for all meteorological subdivisions of India, the recorded monsoon total rainfall data do not include any year with entirely nil rainfall. Hence, the justified probability distributions here are right-skewed and supported over the positive real line; illustrative examples are exponential [todorovic1975stochastic, hazra2018bayesian], gamma [barger1949evaluation, mooley1968application], log-normal [kwaku2007characterization, mandal2015estimation], Weibull [duan1995comparison, lana2017rainfall], and generalized exponential [madi2007bayesian, hazra2025minimum] distributions. However, meteorologists have historically favored the gamma distribution for modeling rainfall data [martinez2019precipitation]. While the data across years are often assumed to be distributed identically in the meteorology literature, it is more appropriate to consider a nonstationary behavior across years in the marginal distribution due to the potential trends in meteorological variables driven by climate change. For this purpose, a gamma regression model is a possible approach that various researchers have explored for different scientific disciplines [nelder1972generalized, amin2020performance].
@@
While the gamma regression can effectively model the nonstationary marginal behavior in rainfall, a fundamental limitation of many traditional univariate models is their failure to incorporate spatial dependency, which is critical for accurate estimation and prediction [diggle1998model]. Copulas [sklar1959distribution, nelsen2006introduction, chen2019copulas] are powerful statistical tools that enable the modeling of complex dependencies between multiple random variables, even extending beyond traditional linear correlation measures [frees1998understanding, omidi2022spatial]. In spatial statistics, copulas facilitate the construction of joint distributions by separately modeling marginal distributions and their spatial dependence structure [sang2010continuous, huser2016non]; this separation allows for greater flexibility in capturing nonlinear and asymmetric dependencies that are often present in spatial data, leading to improved inference and prediction in various applications [graler2011pair, okhrin2017copulae, nazeri2022multivariate, nazeri2023application]. [zhang2012bivariate] attempt to combine the copula theory with the entropy theory for bivariate rainfall and runoff analysis. Further applications of copulas, e.g., Archimedean copulas for meteorological data, are covered in [najjari2012application, zhang2007bivariate]. [mesbahzadeh2019joint] predict temporal trends of precipitation and temperature using a copula under a climate change scenario. In our context of areal data, [musgrove2016hierarchical] propose a copula-based hierarchical model with covariance selection for unbiased estimation of marginal parameters, providing a dependence structure with intuitive conditional and marginal interpretations. They develop a computational framework that permits efficient frequentist inference for their model, even for large datasets. There has been substantial work on copula models with covariate-dependent marginal distributions and dependence structures in recent years. These approaches extend beyond classical static copulas by allowing the copula or marginal parameters to vary with explanatory variables. For example, [patton2006modelling] introduce conditional copulas for time-series applications, while [acar2011dependence] and [klein2016simultaneous] develop regression-type frameworks where covariates enter the copula parameter. Similarly, [abegaz2012semiparametric] and [krupskii2018factor] proposed models with covariate-dependent marginal distributions combined with flexible copula families. Recent contributions include covariate-adjusted vine copulas and spatially varying copula models [czado2019analyzing]. This surge of literature demonstrates a growing interest in models that jointly accommodate marginal heterogeneity and dependence driven by observed or latent covariates.
@@
Several studies have demonstrated the efficacy of using gamma marginal distributions combined with copula models to analyze rainfall patterns. For instance, [khan2020novel] employ gamma marginal distributions within different copula models to capture spatial dependencies in precipitation data. On the other hand, [lee2018multisite] introduces a simulation-based method to enhance the preservation of cross-correlations in multisite precipitation simulations, addressing limitations observed in direct and indirect estimation methods. Here, the author explores a Gaussian copula with identical gamma marginal distributions for data collected at twelve nearby spatial locations; however, the model assumptions are unsuitable for a large geographical domain. Similarly, [van2011copula] utilize copulas with gamma marginal distributions to model the relationship between coarse- and fine-scale rainfall depths. The study in [baxevani2015spatiotemporal] describes precipitation intensity using space and time-dependent gamma distributions. While most of these studies do not assume a gamma regression framework for the marginal distributions, the rest do not focus on annual trends. Besides, the copulas used in these papers are for continuous spatial processes; hence, they are not directly applicable to areal data modeling.
Outside the avenue of copula-based modeling, the two most popular frameworks for areal data modeling are conditional autoregressive (CAR) models [besag1974spatial] and simultaneous autoregressive (SAR) models [cliff1981sar]. While SAR models are used widely in econometrics, regional science, and social sciences, CAR models are popular in Bayesian spatial statistics, epidemiology, and environmental studies. Among numerous uses of the CAR model in the environmental statistics literature, [banerjee2015hierarchical] use hierarchical CAR models for temperature variability analysis, [villarini2010flood] analyze flood risks and rainfall patterns using CAR models, and [cressie2011statistics] use integrated CAR models with spatial dynamic processes for environmental prediction. In most of these models, the data are assumed to be Gaussian, conditioning on the parameters and the latent variables. However, using the CAR model as a copula in the context of environmental statistics or other scientific disciplines is not well-known. Several works have extended areal models beyond the Gaussian framework to accommodate non-Gaussian outcomes such as counts, rates, rainfall indices, or other skewed environmental data. [besag1974spatial] introduce CAR models, which specify the distribution of a spatial variable in a given region conditional on the values in neighboring regions, laying the foundation for modeling spatial dependence in areal data. Subsequent Bayesian hierarchical extensions of CAR models have been developed for disease mapping [best2005comparison, lee2011comparison, jin2005generalized], allowing more flexible modeling of spatially structured risk. More general treatments of hierarchical non-Gaussian areal models can be found in [banerjee2015hierarchical]. Beyond disease mapping, ecology, demography, and environmental sciences applications have used Poisson or Binomial CAR-type models to handle count and rate data [wakefield2007disease, banerjee2015hierarchical]. In recent years, researchers have also developed flexible extensions for overdispersed or skewed data, such as negative binomial CAR models [lawson2018bayesian], zero-inflated areal models [neelon2010bayesian], and generalized additive mixed models with spatial random effects for non-Gaussian responses [wood2017generalized]. Against this backdrop, our contribution differs in two key respects. First, we focus on gamma regression-based marginal distributions tailored to monsoon rainfall totals, linking parameters directly to spatial regions. Second, we combine the marginal distributions with a Gaussian CAR copula that captures spatial dependence across regions. To our knowledge, this combination is novel and has not been applied in the context of large-scale rainfall analysis.
@@
Given the large spatial extent of the meteorological subdivisions of India, with several subdivisions in the western parts of the country receiving very low rainfall and specific subdivisions in the east covering the wettest locations of the globe, the requirement of spatially varying coefficients (SVCs) in gamma regression is obvious [gelfand2003spatial]. Additionally, the regression coefficients are likely to express local homogeneity; hence, we can model the priors for the SVCs within the framework of Gaussian graphical models [banerjee2015hierarchical]. The CAR and intrinsic CAR [ICAR, mollie1996bayesian] priors are popular choices for SVCs. While [cressie2011statistics] discuss ICAR priors in spatio-temporal ecological models, [banerjee2015hierarchical] use ICAR priors for modeling spatial rainfall variability. [fernandez2010bayesian], [lee2015bayesian], and many others use the CAR priors, while the R package CARBayes [lee2017spatio] implements various Bayesian areal data models with CAR priors.
@@
In this paper, we develop a statistical model to analyze monsoon total rainfall data for 34 meteorological subdivisions of mainland India from 1951 to 2014. Given the suitability of the gamma distribution for modeling rainfall data and the possible trend in the marginal distribution due to climate change, we assume a gamma regression model for the marginal distributions, where we consider a logarithmic link for the mean structure, written as a linear function of the years (suitably centered and scaled). We consider a Gaussian CAR copula for modeling the dependence structure, where the adjacency is determined by whether or not two meteorological subdivisions share a boundary. While the usual CAR models exhibit nonstationary marginal distributions due to boundary effects, we appropriately scale the CAR covariance matrix to ensure that the copula conditions are satisfied. We allow varying gamma regression coefficients within a latent Gaussian model framework to account for the natural variation in total rainfall during the monsoon season across different climatic regions. Similar to the likelihood layer, the neighborhood structure of the regions determines the dependence structure in prior layers, where we explore CAR and ICAR priors. Our methodology also effectively imputes missing data by assuming Missing at Random (MAR) [Rubin1976]. We draw Bayesian inference using Markov chain Monte Carlo algorithms, specifically, using a Metropolis-within-Gibbs sampler. To quantify the importance of modeling spatial dependence at the likelihood and prior levels, we compare four model variants: (i) independent likelihood and independent priors, (ii) independent likelihood and CAR/ICAR priors, (iii) CAR-copula likelihood and independent priors, and (iv) CAR-copula likelihood and CAR/ICAR priors. Using various criteria for model comparison, we then compare the performances of the proposed and competing models in terms of model fitting and estimation uncertainties. Applying this method to the Indian areal rainfall dataset, we draw inferences about the model parameters and examine the potential impacts of climate change on rainfall patterns across India. The primary objective of this work is to focus on a specific scientific problem, build a statistical model tailored to that problem, validate the approach through simulation studies, and draw inferences based on the fitted model. In doing so, the contribution does not fall solely within the realm of pure statistical methodology, since we do not aim to provide a fully general modeling solution for areal datasets with spatially varying coefficients; nor is it purely interdisciplinary, as we do not restrict ourselves to applying existing tools without methodological input. Instead, the work lies in between, motivated by an important applied context, while simultaneously contributing to modeling ideas of interest to the statistical community. Specifically, we formulate a Bayesian latent Gaussian model that integrates spatially varying gamma regression with a Gaussian CAR copula, enabling us to model skewed distributions, spatial dependence, and region-specific variability in rainfall. While gamma regression and CAR priors have each been used separately in environmental statistics, to our knowledge, their combination within a copula framework for areal rainfall data has not been previously explored. This contribution is particularly relevant for statisticians seeking flexible models for non-Gaussian, spatially dependent areal data beyond the specific rainfall application. The Indian rainfall dataset serves as a motivating case study to demonstrate the practical utility of the proposed method. Still, the modeling framework applies to various spatial datasets in climate science, epidemiology, and environmental monitoring.
@@
We structure the remainder of this article as follows. Section 2 provides a brief background on gamma regression, conditional autoregressive models, and copulas. In Section 3, we describe the details of the Indian rainfall dataset we analyze and provide an exploratory analysis motivating us to choose specific components in the proposed statistical methodology. Section 4 details the construction of the conditional autoregressive copula for modeling spatial dependence, a gamma regression with SVCs for marginal distributions, and the CAR and ICAR prior structures for the coefficients. In Section 5, we discuss an MCMC sampling strategy employed for parameter estimation. Section 6 discusses a simulation study to validate the correctness of the proposed parameter estimation procedure and to compare the proposed model with simpler alternatives that ignore spatial dependence structures. In Section 7, we apply the proposed and competing methods to the Indian rainfall dataset and discuss the results. Finally, Section 8 concludes with a discussion on key findings, implications for climate studies, and future research directions.
@@ 2 Background
This section provides a brief overview of the key statistical tools required for our methodology and data analysis, including gamma regression for handling nonstationary, skewed, and positive-valued data, conditional autoregressive (CAR) models for areal datasets, and copulas. We also discuss two exploratory tools: the Quantile-Quantile (QQ) plot of uniform-transformed data for marginal model validation and Moran’s I for assessing spatial autocorrelation. Finally, we mention the Deviance Information Criterion (DIC) and the Watanabe-Akaike Information Criterion (WAIC) for Bayesian model comparison.
@@ 2.1 Gamma regression
In a gamma regression model [nelder1972generalized, McCullagh2019glm] for $n$ independent but non-identically distributed positive-valued observations, we assume that the response variable $Y_{i}$ follows
$Y_{i}\overset{\textrm{Indep}}{\sim}\text{Gamma}(\mu_{i},a),~~i=1,\ldots,n,$ (1)
@@
$\lambda_{i}=a/\mu_{i}=a~\exp[-\bm{x}_{i}^{\intercal}\bm{\beta}]=\exp[\log(a)-\bm{x}_{i}^{\intercal}\bm{\beta}].$
Here, the coefficient of variation $\sqrt{\text{Var}(Y_{i})}/\text{E}(Y_{i})=a^{-1/2}$ does not depend on $\mu_{i}$. This property allows us to interpret the shape parameter $a$ as controlling relative dispersion independently of the mean, which will later help us distinguish between regions with similar average rainfall but different variability.
2.2 Conditional Autoregressive (CAR) models
CAR models specify a spatial dependence structure where the conditional distribution of the response variable at each region, given all other areas, depends only on the responses of the neighboring areas [besag1974spatial, cressie1993statistics]. Formally, let $Y_{i}$ denote the response at region $i$, and let $i\sim j$ indicate regions $i$ and $j$ are neighbors. The neighborhood structure is encoded in a spatial weight matrix $\bm{W}$, with its $(i,j)$th element given by
$w_{ij}=\begin{cases}1&\text{if }i\sim j,\\ 0&\text{if }i\nsim j,\end{cases}$ (3)
and $w_{ii}=0$ for all $i$. The number of neighbors of region $i$ is $m_{i}=\sum_{j}w_{ij}$, and we denote by $\bm{M}$ the diagonal matrix with entries $m_{ii}=m_{i}$.
Given a set of regionally indexed observations, the CAR model is defined through the set of full conditional distributions
$Y_{i}\,\big|\,\{Y_{j}:j\neq i\}\;\sim\;\textrm{Normal}\!\left(\rho\sum_{j\in\{i:i\sim j,i\neq j\}}\frac{w_{ij}}{m_{i}}\,Y_{j},\;\frac{\sigma^{2}}{m_{i}}\right),$ (4)
where $\rho$ is a spatial autocorrelation parameter that controls the strength of dependence among the regions, and $\sigma^{2}$ is a marginal variance parameter that controls the overall spatial covariance. Here (4) shows that the conditional mean of $Y_{i}$ is a weighted average of the values of its neighbors, and the conditional variance decreases with the number of neighbors.
By Brook’s lemma [brook1964distinction], the above conditionals correspond to a zero-mean Gaussian Markov random field (GMRF) with precision matrix
$\bm{Q}=\frac{1}{\sigma^{2}}\,(\bm{M}-\rho\bm{W}),$
so that the covariance matrix is $\sigma^{2}(\bm{M}-\rho\bm{W})^{-1}$ whenever this inverse exists. Because the conditional expectations in (4) are expressed entirely in terms of neighboring observations, without any additive constants, the implied joint distribution has a zero mean vector [besag1974spatial, rue2005gaussian, banerjee2015hierarchical].
A proper CAR model arises when the precision matrix $\bm{Q}$ is positive definite, i.e., when $|\rho|<1$ under standard neighborhood choices; this ensures that the joint distribution is a valid multivariate normal distribution. In contrast, when $\rho=1$, the precision matrix is singular, leading to the intrinsic CAR (ICAR) model. The ICAR model does not define a proper joint distribution. Still, it is widely used as an improper prior for latent spatial random effects (e.g., spatially varying coefficients). CAR models thus provide a convenient mechanism for local spatial smoothing, borrowing strength across neighboring areas, and improving predictive accuracy and inference.